If a+b+c=0 and |a|=3,|b|=5,|c|=7, then angle between vectors a and b is??

Grade:12th pass

1 Answers

Ashik Raveendran
5 years ago
a + b +c =0
then, a + b =-c
taking modulus on both sides ,\left |a+b \left | = \left | -c \right |
\Rightarrow \left | a+b \right |^{2}=\left | -c^{} \right |^{2}
a^{2}+2\vec{a}\vec{b}+b^{2}=c^{2}
abcos\theta =c^{2}-a^{2}-b^{2}
\Rightarrow cos\theta =c^{2}-a^{2}-b^{2}/2ab
substituting the values we get ,cos\theta =1/2
\Rightarrow \theta =60^{o}
All the bes